OpenStudy (anonymous):

the ratio of 8:5= (?)% please break it down in the lowest terms

OpenStudy (anonymous):

62.5%

OpenStudy (anonymous):

the answer is 160% but how

OpenStudy (anonymous):

oh, then it's backwards from what I'm used to

OpenStudy (anonymous):

anyway, its just 8/5

OpenStudy (anonymous):

nevermind, I'm the one thats backwards

OpenStudy (anonymous):

i have multiple choice here and they r 75%, 150%, 175% and 125% but neither one of those r correct

OpenStudy (anonymous):

the correct one is suppose to b 160%

OpenStudy (anonymous):

i don't know what to tell you, but you are right

OpenStudy (anonymous):

8/5*x/100=160=x

OpenStudy (anonymous):

5/800=160

OpenStudy (across):

\[\frac{x}{8}=\frac{100}{5}\]Solve for x.